### v1

@@ Abstract
In this paper we investigate the spatio-temporal dynamics of obesity rates across Italian regions from 2010 to 2022, aiming to identify spatial and temporal trends and assess potential heterogeneities. We implement a Bayesian hierarchical Beta regression model to analyze regional obesity rates, integrating spatial and temporal random effects, alongside gender and various exogenous predictors. The model leverages the Stochastic Search Variable Selection technique to identify significant predictors supported by the data. The analysis reveals both regional heterogeneity and dependence in obesity rates over the study period, emphasizing the importance of considering gender and spatial correlation in explaining its dynamics over time. In fact, the inclusion of structured spatial and temporal random effects captures the complexities of regional variations over time. These random effects, along with gender, emerge as the primary determinants of obesity prevalence across Italian regions, while the role of exogenous covariates is found to be minimal at the regional level. While socioeconomic and lifestyle factors remain fundamental at a micro-level, the findings demonstrate that the integration of spatial and temporal structures is critical for capturing macro-level obesity variations.
@@ 1 Introduction
Obesity is recognized as a major public health problem in Europe, where it has reached alarming levels, posing substantial health, social, and economic burdens on European societies. In 2014, one over six adults were considered obese according to the European Health Interview Survey (2016) [^1] Eurostat news release, 203/2016, 20 October 2016, https://ec.europa.eu/eurostat/documents/2995521/7700898/3-20102016-BP-EN.pdf/c26b037b-d5f3-4c05-89c1-00bf0b98d646. Moreover, projections suggest an increasing average obesity trend (Janssen et al.,, 2020).
@@
Even if the problem is common among countries, such a phenomenon is not homogeneous: overweight and obesity vary substantially between and within countries. For instance, compared to other European countries, Italy shows relatively lower obesity rates for women (OECD,, 2022). According to the Eurostat data (Eurostat,, 2022), the second lowest proportion of obese people in 2019 was recorded in Italy (slightly above 11%), when nearby countries like Spain, France, Germany, and Austria are all above 14 %, and the European mean is around 17%. Despite this, the situation is still problematic and needs appropriate attention and intervention.
@@
Moreover, national averages hide substantial heterogeneity within the countries, both in space and by gender, through time. Focusing on Italy, between northern and southern regions, there is a notable and time consistent difference in overweight and obesity rates. For example, taking the best and worst regions, in the period considered by this study (2010-2022), the female obesity rate (male) is 7.50% (9.40%) in Trentino-Alto Adige and 12.7% (14.4%) in Molise. While considering overweight for the same regions, the percentages are 23.5% (40.3%) for Trentino-Alto Adige and 31% (46%) for Molise. It is well-known that across Italian regions, there are strong structural differences in general culture, social norms, and economic development (for a general discussion, see, for instance, De Blasio and Nuzzo,, 2009). The diversity observed among Italian regions may influence obesity and related phenomena (Gallus et al.,, 2013). For this reason, proposing an analysis that takes into account heterogeneity is essential (Brunello and Labartino,, 2014). In a related direction, flexible models - such as varying coefficient models - have also been proposed to explore time-varying and spatially-varying effects on public health outcomes, including obesity, particularly using large-scale survey data such as the Behavioral Risk Factor Surveillance System (Assaf et al.,, 2015). Furthermore, gender aspects should also be considered, given potential physical and genetic differences (Varì,, 2016) as well as cultural ones (Barra and Barone,, 2021).
@@
In this paper, we analyse longitudinal areal data consisting of yearly regional obesity rates from 2010 to 2022. To model these rates, expressed as proportions bounded between 0 and 1, we employ a Beta regression model within a Bayesian hierarchical framework. This approach provides a flexible structure for modelling rate data, enabling the inclusion of covariates as well as unobserved spatial and temporal heterogeneity through structured random effects. The Beta regression framework has been extensively studied over the past two decades. For instance, Simas et al., (2010) extend Beta regression by modelling both the mean parameter and the precision parameter as linear combinations of covariates, linked via appropriate functions. Similarly, Da Silva and De Oliveira Lima, (2017) incorporate Beta regression into a geographically weighted regression framework, adapting the model to account for spatial variation. Furthermore, Figueroa-Zúñiga et al., (2013) extend Beta regression by embedding it within a Bayesian hierarchical framework, enabling the inclusion of random effects. Our Beta regression model follows the Bayesian hierarchical framework proposed by Figueroa-Zúñiga et al., (2013), but introduces key differences in the specification of random effects and in the approach used for variable selection. These modifications are designed to address the specific challenges of modelling spatio-temporal structures in areal data and to facilitate effective selection from a large set of covariates. Bayesian Beta regression models with spatio-temporal random effects have been explored in the literature, such as in the work of Lagos-Alvarez et al., (2017), which focuses on geostatistical data rather than areal units. However, to the best of our knowledge, no existing Bayesian Beta regression model for areal data simultaneously incorporates spatio-temporal random effects, group-specific intercepts, and Bayesian variable selection. We mention that other approaches to deal with rate data have been studied. For example, compositional data analysis has been proposed to jointly forecast multiple weight categories (e.g., underweight, overweight, obese) using log-ratio transformations (Mills,, 2008).
@@
To account for spatial and temporal correlation, we incorporate an additive formulation for the random effects in the linear predictor of the Beta regression model. Specifically, we assume an Intrinsic Conditional Auto-Regressive (ICAR) prior for the spatially structured random effect, exploiting its ability to capture structured spatial dependence, which is particularly relevant for regional-level data where geographic proximity strongly influences observed outcomes (Banerjee et al.,, 2014; Haining and Li,, 2020). The temporal random effect is modelled as a first-order auto-regressive process (AR(1)). Furthermore, in order to capture potential group differences across gender, we include gender-specific intercept terms in the model. This allows to assess the influence of gender alongside spatial and temporal effects. In this way, our approach jointly considers space, time, and gender influences on regional obesity rates, all of which have been shown to be highly relevant (Grimaccia and Rota,, 2025). The additive combination of the spatial and temporal random effect offers a parsimonious and interpretable framework, making it a suitable modelling strategy (Banerjee et al.,, 2014). Although more structured interactions between spatial and temporal components could be incorporated (for a review, see Knorr-Held,, 2000), we follow the approach of Waller et al., (1997) and Berloco et al., (2023), arguing that prioritizing simplicity and interpretability over added flexible and complexity aligns with the primary objective of this study: to effectively model regional-level patterns in obesity rates while focusing on structured spatio-temporal dynamics.
@@
With the objective of finding a parsimonious representation of the observed data using a statistical model that is also capable of accurate prediction, we adopt the Stochastic Search Variable Selection (SSVS) approach (George and McCulloch,, 1993, George and McCulloch,, 1997). This method is a Bayesian regularization technique that has proven effective in addressing both prediction and variable selection problems (see Rockova et al.,, 2012, for a detailed discussion). SSVS is particularly attractive because it is well-suited for high-dimensional modelling, provides valid standard errors, and enables the simultaneous estimation of regression coefficients and complexity parameters through computationally efficient Markov chain Monte Carlo (MCMC) techniques.
@@
The paper is structured as follows. In Section 2, we detail the construction of the dataset, which was assembled from multiple official sources. In Section 3 we outline the Bayesian hierarchical spatio-temporal Beta regression model (Section 3.1), providing details about prior assumptions (Section 3.2) for the structured spatial and temporal random effects, alongside the gender-specific intercepts. Section 3.3 and 3.4 discuss the SSVS prior specification for the fixed effects and some computation details, respectively. We then present the results of our analysis in Section 4 and some final considerations in Section 5. Additional material is provided in the appendices: Appendix 10 contains the table describing the covariates used in the analysis; Appendix 11 reports additional posterior summary statistics not included in the main text for brevity; and Appendix 12 presents the results of the comparison of some variants of the main model.
@@ 2 Data
The dataset analysed in this study comprises spatio-temporal areal data, which can be conceptualized as a set of time series associated with each spatial unit (i.e., region). The dataset is organized as a longitudinal dataset, collecting information on the 20 Italian regions from 2010 to 2022 (13 years). The Italian regions are divided, according to the NUTS 1 classification [^2] NUTS classification: https://ec.europa.eu/eurostat/web/nuts into 5 different geographical zones: North West (NO), North East (NE), Centre (C), South (S) and Islands (SI). In Table 1 we report the regions and the corresponding geographical zone.
Zone Regions
North East (NE) Trentino-Alto Adige, Veneto, Friuli-Venezia Giulia
Centre (C) Toscana, Umbria, Marche, Lazio
South (S) Abruzzo, Molise, Campania, Puglia, Basilicata, Calabria
Islands (SI) Sicilia, Sardegna
[…]
@@
The dataset contains aggregated data at the regional level (the NUTS 2 level) with variables belonging to four groups: socio-demography, mortality, lifestyle, and general health-related information (see Table 7 in Appendix 10). We assembled the dataset by extracting information from the Health For All platform (HFA [^3] HFA: https://www.istat.it/it/archivio/14562), the Benessere Equo e Sostenibile (BES [^4] BES: https://www.istat.it/en/well-being-and-sustainability/the-measurement-of-well-being/bes-report) report, the Eurostat databases [^5] Eurostat: https://ec.europa.eu/eurostat/data/database and the archive of the Ministry of Economy and Finance (MEF [^6] MEF archive: www.rgs.mef.gov.it/VERSIONE-I/archivio).
@@
The dependent variable is the obesity rate measured yearly for each region in the time window considered separately by gender (source: HFA). The definition of obesity follows the WHO classification [^7] WHO Obesity definition: https://www.who.int/news-room/fact-sheets/detail/obesity-and-overweight, i.e., having a Body Mass Index (BMI) greater than or equal to 30 defines an obese individual.
@@
For the mortality rate variables in Group 2 of Table 7, which exhibited high multicollinearity and strong linear relationships, we applied a principal component analysis (PCA). These variables represent cause-specific mortality rates, calculated as the number of deaths due to a given cause per 10,000 individuals within the same age and gender group. Among the nine principal components generated, we retained the first two, which together explained more than 70% of the total variance. The first principal component (PC1) primarily captures diseases related to cardiovascular issues and diabetes, while the second (PC2) reflects factors more associated with stress and mental disorders.
@@ 3.1 Beta regression model
Let $Y_{its}$ denote the obesity rate in region $i$ $(i=1,\dots,20)$, year $t$ $(t=1,\dots,13)$ and gender $s$ $(s=1\text{ for females and}$ $s=2\text{ for males})$. We specify the following Beta regression model (Ferrari and Cribari-Neto,, 2004):
@@
With this formulation we have that $\mu_{its}=E(Y_{its})\in(0,1)$ and $\text{Var}(Y_{its})=\left(\mu_{its}(1-\mu_{its})\right)/(1+\phi)$. It follows that $\mu_{its}$ is the mean obesity rate for region $i$ at year $t$ and for gender $s$, and the parameter $\phi$ can be interpreted as a constant precision parameter, i.e., given $\mu_{its}$, the larger the value of $\phi$, the smaller the variance of $Y_{its}$. The structured space random effect is represented by $\bm{\psi}=(\psi_{1},\psi_{2},\ldots,\psi_{20})^{T}$, while the structured time random effect by $\bm{\alpha}\ =(\alpha_{1},\alpha_{2},\ldots,\alpha_{13})^{T}$. The global intercept term is $\beta_{0}$, while the gender-specific intercepts are introduced through the dummy variable $d_{s}$ which takes the value 0 for females and 1 for males. Following Berloco et al., (2023), we adopt an additive specification for the spatial and temporal effects. More details about the associated prior specification are provided in Section 3.2. The fixed effects are introduced in the model through the vector of $p=32$ exogenous covariates (see Table 7 in Appendix 10), denoted by $\bm{x}_{its}=(x_{its1},x_{its2},\ldots,x_{itsp})^{T}$, and the corresponding vector of coefficients $\bm{\beta}=(\beta_{1},\beta_{2},\ldots,\beta_{p})^{T}$.
The Beta parametrization adopted in our model offers distinct advantages. First, by expressing the distribution in terms of its mean $\mu_{its}$ and precision $\phi$, the model facilitates a more intuitive interpretation. Second, this formulation simplifies the specification of prior distributions within the Bayesian framework. Prior knowledge can be naturally incorporated through $\mu_{its}$ via the linear predictor. Simultaneously, the precision parameter $\phi$ can be assigned a weakly informative prior, allowing the model to flexibly adapt to the data’s inherent variability (Ferrari and Cribari-Neto,, 2004).

### v2

@@ Abstract
In this paper, we investigate the spatio-temporal dynamics of obesity across Italian regions from 2010 to 2022, introducing a novel way to analyse aggregated regional obesity data. In particular, we implement a Bayesian hierarchical Beta regression model to analyse regional obesity rates, integrating spatial and temporal random effects, alongside gender and several exogenous predictors. The model leverages the Stochastic Search Variable Selection technique to identify significant predictors supported by the data. The analysis reveals both gender and regional heterogeneity in obesity rates over the study period. Structured spatial and temporal random effects, along with gender, emerge as the primary determinants of obesity prevalence across Italian regions, while the role of exogenous covariates is found to be minimal at the regional level, once structured spatial and temporal effects and gender are accounted for. While socioeconomic and lifestyle factors remain fundamental at a micro-level, the findings suggest that the integration of spatial and temporal structures is crucial for capturing macro-level obesity variations.
@@ 1 Introduction
Obesity is recognized as a major public health problem in Europe, where it has reached alarming levels, posing substantial health, social, and economic burdens on European societies. In 2014, one in six adults were considered obese according to the European Health Interview Survey (2016) [^1] Eurostat news release, 203/2016, 20 October 2016, https://ec.europa.eu/eurostat/documents/2995521/7700898/3-20102016-BP-EN.pdf/c26b037b-d5f3-4c05-89c1-00bf0b98d646. Moreover, projections suggest an increasing average obesity trend [22].
@@
Even if the problem is common among countries, such a phenomenon is not homogeneous: overweight and obesity vary substantially between and within countries. For instance, compared to other European countries, Italy shows relatively lower obesity rates for women [29]. According to the Eurostat data [11], the second lowest proportion of obese people in 2019 was recorded in Italy (slightly above 11%), when nearby countries like Spain, France, Germany, and Austria are all above 14 %, and the European mean is around 17%. Despite this, the situation is still problematic and needs appropriate attention and intervention.
@@
Moreover, national averages hide substantial heterogeneity within the countries, both in space and by gender, through time. Focusing on Italy, between northern and southern regions, there is a notable and time consistent difference in overweight and obesity rates. For example, taking the best and worst regions in the period considered by this study (2010-2022), our data show that the female (male) obesity rate is 7.50% (9.40%) in Trentino-Alto Adige and 12.7% (14.4%) in Molise. While considering overweight for the same regions, the percentages are 23.5% (40.3%) for Trentino-Alto Adige and 31% (46%) for Molise. It is well-known that across Italian regions, there are strong structural differences in general culture, social norms, and economic development (for a general discussion, see, for instance, 10). This diversity among Italian regions may influence obesity and related phenomena [14]. For this reason, proposing an analysis that takes into account heterogeneity is essential [7]. Flexible models, such as varying coefficient models, have also been proposed in [1] to explore spatially and temporally varying effects on public health outcomes, including obesity, using large-scale survey data. Their findings show that this framework can capture dynamic changes over space and time in the effects of key risk factors, providing insight into the evolution of obesity across population subgroups. Their focus, however, differs from ours: while they concentrate on time and subgroup-specific variation in covariate effects, we are mainly interested in studying and interpreting the spatio-temporal pattern of obesity rates after accounting for covariates. Along with space and time, gender aspects should also be considered, given potential physical and genetic differences [38] as well as cultural ones [3].
@@
In this paper, we analyse longitudinal areal data consisting of yearly regional obesity rates from 2010 to 2022, which were assembled through a systematic process of data collection, validation, and harmonization across multiple institutional sources. To model these rates, expressed as proportions bounded between 0 and 1, we employ a Beta regression model within a Bayesian hierarchical framework, using a mean–precision parameterization. This approach provides a flexible structure for modelling rate data, enabling the inclusion of covariates as well as unobserved spatial and temporal heterogeneity through structured random effects. The Beta regression framework has been extensively studied over the past two decades. For instance, Simas et al. [34] extend Beta regression by modelling both the mean parameter and the precision parameter as linear combinations of covariates, linked via appropriate functions. Similarly, Da Silva and De Oliveira Lima [9] incorporate Beta regression into a geographically weighted regression framework, adapting the model to account for spatial variation. Furthermore, Figueroa-Zúñiga et al. [13] extend Beta regression by embedding it within a Bayesian hierarchical framework, enabling the inclusion of random effects. The Beta regression model we propose builds on the Bayesian hierarchical framework of Figueroa-Zúñiga et al. [13], while introducing important modifications to the random-effects specification and the variable selection strategy to address the specific challenges of modelling spatio-temporal structures in areal data and selecting from a large set of covariates. Bayesian Beta regression models with spatial dependence have been developed mainly in geostatistical frameworks, where Gaussian spatial random fields enter in the linear predictor to capture distance-based correlation [26, 30]. For areal data, spatial dependence is more naturally encoded through neighbourhood structures, as in Cepeda-Cuervo and Núñez-Antón [8], where a spatial autoregressive Beta regression with contiguity-based lags in both mean and dispersion is used.
@@
To account for spatial and temporal correlation, we incorporate in the linear predictor of the Beta regression model an additive formulation for the random effects. Specifically, we assume an Intrinsic Conditional Auto-Regressive (ICAR) prior for the spatially structured random effect, exploiting its ability to capture structured spatial dependence, which is particularly relevant for regional-level data where geographic proximity strongly influences observed outcomes [2, 20]. The temporal random effect is modelled as a first-order autoregressive process (AR(1)). Furthermore, in order to capture potential group differences across gender, we include gender-specific intercept terms in the model. In this way, our approach jointly accounts for space, time, and gender influences on regional obesity rates, all of which have been shown to be highly relevant [19]. The additive combination of the spatial and temporal random effect offers a parsimonious and interpretable framework, making it a suitable modelling strategy [2]. Although more structured interactions between spatial and temporal components could be incorporated [for a review, see 25], we follow the approach of Waller et al. [40] and Berloco et al. [4], arguing that prioritizing simplicity and interpretability over added flexible and complexity aligns with the primary objective of this study: to effectively model regional-level patterns in obesity rates while focusing on structured spatio-temporal dynamics.
@@
With the objective of finding a parsimonious representation of the observed data using a statistical model that is also capable of accurate prediction, we adopt the Stochastic Search Variable Selection (SSVS) approach (17, 18). This method is a Bayesian regularization technique that has proven effective in addressing both prediction and variable selection problems [see 32, for a detailed discussion]. SSVS is particularly attractive because it is well-suited for high-dimensional modelling, provides valid standard errors, and enables the simultaneous estimation of regression coefficients and complexity parameters through computationally efficient Markov chain Monte Carlo (MCMC) techniques.
Although the individual components of our proposed approach, namely Bayesian Beta regression, ICAR spatial effects, AR(1) temporal dynamics, group-specific intercepts, and SSVS, are all well established in the literature, their integration within a single coherent Bayesian hierarchical framework for modeling regional obesity rates is, to our knowledge, novel. Our contribution lies in combining these elements into an areal spatio-temporal model that jointly accounts for spatial and temporal heterogeneity, gender-specific differences, and covariate uncertainty, thereby providing a flexible and comprehensive framework for analysing regional obesity dynamics.
Beyond this methodological integration, we construct a harmonized regional panel dataset on obesity and related covariates for Italy through a systematic process of data collection, validation, and reconciliation across multiple institutional sources. This integrated modelling and data effort enables robust inference and improved prediction of obesity prevalence at the regional level, delivering evidence directly relevant for place-based public health policies.
@@
The paper is structured as follows. In Section 2, we detail the construction of the dataset, which was assembled from multiple official sources. In Section 3 we outline the Bayesian hierarchical spatio-temporal Beta regression model (Section 3.1), providing details about prior assumptions (Section 3.2) for the structured spatial and temporal random effects, alongside the gender-specific intercepts. Section 3.3 and 3.4 discuss the SSVS prior specification for the fixed effects and some computation details, respectively. We then present the results of our analysis in Section 4 and some final considerations in Section 5. Additional material, including descriptive statistics, posterior summaries, predictive comparisons, geographical cross-validation, and robustness diagnostics, is provided in the appendices.
@@ 2 Data
The dataset analysed in this study comprises spatio-temporal areal data, which can be conceptualized as a set of time series associated with each spatial unit (i.e., region). The dataset is organized as a longitudinal dataset, collecting information on the 20 Italian regions from 2010 to 2022 (13 years). The Italian regions are divided, according to the NUTS 1 classification [^2] NUTS classification: https://ec.europa.eu/eurostat/web/nuts into 5 different geographical zones: North West (NO), North East (NE), Centre (C), South (S) and Islands (SI). In Table 6 (Appendix B) we report the regions and the corresponding geographical zone.
In our analysis, a “weights matrix” $\bm{W}$ is needed to model spatial dependence through the ICAR prior (see Section 3). This matrix encodes the neighborhood relationships between regions, and due to its binary nature, a challenge arises in defining neighbors for the two main Italian islands, Sicilia and Sardegna. We treat Sicilia as a neighbor of Calabria, reflecting their geographic proximity across the Strait of Messina. In contrast, Sardegna is modeled as isolated, given its substantial spatial separation from all other regions. This specification was selected after comparing the values of Moran’s index [see 2] under alternative neighborhood definitions for the two islands [see 19]
@@
The dataset contains aggregated data at the regional level (NUTS 2) with variables grouped into four categories: socio-demography, mortality, lifestyle, and general health-related information (see Table 7 in Appendix B). We assembled the dataset by extracting information from several official sources: the Health For All platform (HFA [^3] HFA: https://www.istat.it/it/archivio/14562), the Benessere Equo e Sostenibile (BES [^4] BES: https://www.istat.it/en/well-being-and-sustainability/the-measurement-of-well-being/bes-report) report, Eurostat databases [^5] Eurostat: https://ec.europa.eu/eurostat/data/database and the archive of the Ministry of Economy and Finance (MEF [^6] MEF archive: www.rgs.mef.gov.it/VERSIONE-I/archivio).
Figure 1: Obesity rate trend from 2010 to 2022 by geographic area and gender.
@@
The dependent variable is the obesity rate measured yearly for each region and considered separately by gender (source: HFA). We display in Figure 1 the obesity trends by geographic area and gender for the time window we consider. The definition of obesity follows the WHO classification [^7] WHO Obesity definition: https://www.who.int/news-room/fact-sheets/detail/obesity-and-overweight, i.e., having a Body Mass Index (BMI) greater than or equal to 30 defines an obese individual. In the HFA platform, obesity rates are official ISTAT sample-based estimates derived from multipurpose household surveys.
@@
For the mortality rate variables in Group 2 of Table 7, which exhibited high multicollinearity and strong linear relationships, we applied a principal component analysis (PCA). These variables represent cause-specific mortality rates, calculated as the number of deaths due to a given cause per 10,000 individuals within the same age and gender group. Among the nine principal components generated, we retained the first two, which together explained more than 70% of the total variance. The first principal component (PC1) primarily captures diseases related to cardiovascular issues and diabetes, while the second (PC2) reflects factors more associated with stress and mental disorders. Summing up, the final dataset contains $p=32$ covariates (see Table 7 in Appendix B for more details).
@@ 3.1 Beta regression model
Let $Y_{its}$ denote the obesity rate in region $i$ $(i=1,\dots,20)$, year $t$ $(t=1,\dots,13)$ and gender $s$ $(s=1\text{ for females and}$ $s=2\text{ for males})$. We specify the following Beta regression model [12]:
@@
With this formulation we have that $\mu_{its}=E(Y_{its})\in(0,1)$ and $\text{Var}(Y_{its})=\left(\mu_{its}(1-\mu_{its})\right)/(1+\phi)$. It follows that $\mu_{its}$ is the mean obesity rate for region $i$ at year $t$ and for gender $s$, and the parameter $\phi$ can be interpreted as a precision parameter, i.e., given $\mu_{its}$, the larger the value of $\phi$, the smaller the variance of $Y_{its}$. The structured space random effect is represented by $\bm{\psi}=(\psi_{1},\psi_{2},\ldots,\psi_{20})^{T}$, while the structured time random effect by $\bm{\alpha}\ =(\alpha_{1},\alpha_{2},\ldots,\alpha_{13})^{T}$. The global intercept term is $\beta_{0}$, while the gender-specific intercepts are introduced through the dummy variable $d_{s}$, which takes the value 0 for females and 1 for males.
Following Berloco et al. [4], we adopt an additive specification for the spatial and temporal effects. More details about the associated prior specification are provided in Section 3.2. We denote the covariate vector with $\bm{x}_{its}=(x_{its1},x_{its2},\ldots,x_{itsp})^{T}$ and the corresponding fixed effect coefficients with $\bm{\beta}=(\beta_{1},\beta_{2},\ldots,\beta_{p})^{T}$.
The Beta parametrization adopted in our model offers distinct advantages. First, by expressing the distribution in terms of its mean $\mu_{its}$ and precision $\phi$, the model facilitates a more intuitive interpretation. Second, this formulation simplifies the specification of prior distributions within the Bayesian framework. Prior knowledge can be naturally incorporated through $\mu_{its}$ via the linear predictor. Simultaneously, the precision parameter $\phi$ can be assigned a weakly informative prior, allowing the model to flexibly adapt to the data’s inherent variability [12].